A concentration cell was constructed by immersing two silver electrodes in 0.05 M and 1M AgNO3 Solutions. Give the cell representation write the cell reaction and calculate the emf of the cell


1
Expert's answer
2021-01-18T01:12:20-0500

The cell representation: Ag1+ I Ag0 I Ag+1 I Ag0


The cell reactions: A(-) : Ag0 = Ag+1 + 1e

C(+) :Ag1+ +1e = Ag0

To find the EMF of a concentration element, you must use the following formula

"E = \frac{RT}{F}Xln\frac{C1}{C2}"

c1>c2

Therefore,E = 0.08 V